# GeoDANO: Geometric VLM with Domain Agnostic Vision Encoder

### v1

@@ Abstract
We introduce GeoDANO, a geometric vision-language model (VLM) with a domain-agnostic vision encoder, for solving plane geometry problems. Although VLMs have been employed for solving geometry problems, their ability to recognize geometric features remains insufficiently analyzed. To address this gap, we propose a benchmark that evaluates the recognition of visual geometric features, including primitives such as dots and lines, and relations such as orthogonality. Our preliminary study shows that vision encoders often used in general-purpose VLMs, e.g., OpenCLIP, fail to detect these features and struggle to generalize across domains. We develop GeoCLIP, a CLIP-based model trained on synthetic geometric diagram–caption pairs to overcome the limitation. Benchmark results show that GeoCLIP outperforms existing vision encoders in recognizing geometric features. We then propose our VLM, GeoDANO, which augments GeoCLIP with a domain adaptation strategy for unseen diagram styles. GeoDANO outperforms specialized methods for plane geometry problems and GPT-4o on MathVerse.
@@
Seunghyuk Cho1, Zhenyue Qin3, Yang Liu3, Youngbin Choi1, Seungbeom Lee1, Dongwoo Kim1,2 1Graduate School of Artificial Intelligence, POSTECH, 2Department of Computer Science and Engineering, POSTECH, 3Independent Researcher Correspondence to: Dongwoo Kim <dongwoo.kim@postech.ac.kr>
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in automated math problem solving, particularly through code-generation capabilities integrated with proof assistants (Moura and Ullrich, 2021; Nipkow et al., 2002; Chen et al., 2023; Wu et al., 2022; Hendrycks et al., 2021). Although LLMs excel at generating solution steps and correct answers in algebra and calculus (Zhou et al., 2024), their unimodal nature limits performance in plane geometry, where solution depends on both diagram and text (Zhou et al., 2024).
Specialized vision-language models (VLMs) have accordingly been developed for plane geometry problem solving (PGPS) (Chen et al., 2021, 2022; Lu et al., 2021; Zhang et al., 2023; Zhang and Moshfeghi, 2024; Li et al., 2024b; Xia et al., 2024). Yet, it remains unclear whether these models genuinely leverage diagrams or rely almost exclusively on textual features. This ambiguity arises because existing PGPS datasets typically embed sufficient geometric details within problem statements, potentially making the vision encoder unnecessary (Zhang and Moshfeghi, 2024). Fig. 1 illustrates example questions from GeoQA and PGPS9K, where solutions can be derived without referencing the diagrams.
@@
Figure 1: Examples of diagram-caption pairs and their solution steps written in formal languages from GeoQA and PGPS9k datasets. In the problem description, the visual geometric premises and numerical variables are highlighted in green and red, respectively. A significant difference in the style of the diagram and formal language can be observable.
We propose a new benchmark created via a synthetic data engine, which systematically evaluates the ability of VLM vision encoders to recognize geometric premises. Our empirical findings reveal that previously suggested self-supervised learning (SSL) approaches, e.g., vector quantized variataional auto-encoder (VQ-VAE) (Liang et al., 2023) and masked auto-encoder (MAE) (Ning et al., 2023; Xia et al., 2024), and widely adopted encoders, e.g., OpenCLIP (Radford et al., 2021) and DinoV2 (Oquab et al., 2024), struggle to detect geometric features such as perpendicularity and degrees.
To this end, we propose GeoCLIP, a model pre-trained on a large corpus of synthetic diagram–caption pairs. By varying diagram styles (e.g., color, font size, resolution, line width), GeoCLIP learns robust geometric representations and outperforms prior SSL-based methods on our benchmark. Building on GeoCLIP, we introduce a few-shot domain adaptation technique that efficiently transfers the recognition ability to real-world diagrams. We further combine this domain-adapted GeoCLIP with an LLM, forming a domain-agnostic VLM for solving PGPS tasks in MathVerse (Zhang et al., 2024a).
In our experiments on MathVerse (Zhang et al., 2024a), which encompasses diverse plane geometry tasks and diagram styles, our VLM with a domain-adapted GeoCLIP consistently outperforms both task-specific PGPS models and generalist VLMs. Ablation studies confirm the effectiveness of our domain adaptation strategy, showing improvements in optical character recognition (OCR)-based tasks and robust diagram embeddings across different styles.
We summarize the contributions as follows: We propose a novel benchmark for systematically assessing how well vision encoders recognize geometric premises in plane geometry diagrams (Section 3); We introduce GeoCLIP, a vision encoder capable of accurately detecting visual geometric premises (Section 4.1), and a few-shot domain adaptation technique that efficiently transfers this capability across different diagram styles (Section 4.2); We show that our VLM, incorporating domain-adapted GeoCLIP, surpasses existing specialized PGPS VLMs and generalist VLMs on the MathVerse benchmark (Section 5.2) and effectively interprets diverse diagram styles (Section 5.3).
@@ 2 Related Work
In this section, we summarize the studies related to the benchmarks proposed to evaluate plane geometry problem solving (PGPS), the models trained for PGPS, and the contrastive learning methods used to enhance PGPS performance.
@@ 2.1 PGPS benchmarks
However, as recent work by Zhang et al. (2024a) shows, the problem description contains too much information such that the model produces a valid solution program without having the diagram information as shown in Fig. 1. MathVerse (Zhang et al., 2024a) introduces modifications to existing PGPS benchmarks by directly encoding the geometric properties and relations into the diagrams. Therefore, it is impossible to produce a valid solution without recognizing the necessary information from diagrams. Despite the effort, it is still unclear to what extent the vision encoder recognizes the geometric conditions in a diagram as models are evaluated in an end-to-end fashion.
@@ 2.2 Program generation based PGPS
Another line of research typically employs vision-language models (VLMs), where a VLM comprises a vision encoder and a language model (Zhang et al., 2023; Chen et al., 2021; Cao and Xiao, 2022; Ning et al., 2023; Chen et al., 2022; Liang et al., 2023; Xia et al., 2024; Li et al., 2024b). The vision encoder produces a visual embedding from the diagram, and the language model then generates solution steps in an autoregressive manner, conditioned on the textual description and the visual embedding. While the VLMs apply to various diagram formats, the visual geometric premises perception of the VLMs remains underexplored due to the abundance of textual information in existing benchmarks. Moreover, the VLMs are often fine-tuned and tested on a single benchmark, leaving their domain generalization capabilities across different diagram styles unexamined.
@@ 2.3 Contrastive learning in PGPS
Contrastive learning is applied in diverse domains such as computer vision (Schroff et al., 2015) and natural language processing (Gao et al., 2021). In the context of PGPS, contrastive learning is employed to address domain-specific challenges. GeoX (Xia et al., 2024) applies contrastive learning to the adapter layer of the VLM to enhance formal language comprehension. Other approaches train the vision encoder itself using the contrastive language-image pre-training (CLIP) (Radford et al., 2021) objective: LANS (Li et al., 2024b) aligns patch embeddings from a vision Transformer (ViT) with text token embeddings if they describe the same point, and MAVIS (Zhang et al., 2024b) employs diagram–caption pairs generated by a synthetic engine for CLIP. In this work, we examine how CLIP with varied caption styles influences the visual geometric premises recognition of the vision encoder. In addition, a contrastive learning framework is introduced to strengthen robustness against domain shifts in the styles of diagrams.
3 Visual Geometric Premises Recognition Benchmark for Vision Encoders
@@ 3.1 Benchmark preparation
  • Concyclic: A circle and four points are given. The task is to identify how many of those points lie on the circle.
  • TwoLines: Two lines, AB and BC, are given alongside other geometric objects. The task is to determine whether AB and BC are perpendicular, collinear, or neither.
@@
An example of each task is provided in Fig. 2.
@@
Our benchmark is built on top of AlphaGeometry (Trinh et al., 2024), which is designed to solve IMO-style plane geometry problems. The program provides useful functions such as formal language describing plane diagrams. The language predefines a set of geometric premises listed in Table 5, including all necessary properties to define our benchmark tasks. In addition, once a diagram description is given in formal language, the program renders a corresponding diagram with varying fonts, colors, widths, orientations, and resolutions, allowing us to have diagrams with diverse styles often observed in a real-world scenario.
We create question-and-answer pairs based on AlphaGeometry. To sample a diverse set of question-and-answers, we first establish a foundational geometric structure corresponding to the key problem of the task and then repeatedly add new points or lines with randomly selected geometric relationships to the existing diagram with the help of the formal language. The pseudo-code for the random question generation is presented in Algorithm 1. For each task, we generate 50,000, 10,000, and 10,000 question-and-answer pairs for training, validation, and testing, respectively.
@@ 3.2 Results
SSL Jigsaw 86.11 63.85 49.98 61.88 11.44
@@
GeoCLIP GeoCLIP (F $\times$) 99.52 98.61 88.33 86.76 65.68
@@
With the proposed benchmark, we evaluate four widely adopted vision encoders for the open-sourced VLMs: OpenCLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), DinoV2 (Oquab et al., 2024), and ConvNeXT (Liu et al., 2022).
To evaluate the vision encoder, we adopt a linear probing approach. Specifically, we add a linear layer on top of each encoder as a prediction head and train the linear layer from scratch while freezing the parameters of the vision encoder. We use a training set to train the prediction head and report the test accuracy with the best validation performance. The details for the hyper-parameters are described in Section B.1.
As shown in Table 1, many existing vision encoders well recognize the shape of objects but fail to recognize the correct angle between two lines. The encoders also show some difficulties in recognizing the shape of a square and the relationship between two lines. Although the result may seem satisfactory at a glance, these errors will propagate to the downstream tasks when combined with LLMs. Hence, it is important to improve the recognition performance of the vision encoder further.
4 Improving the Vision Encoder Geometric Premises Recognition
In this section, we first propose GeoCLIP, a new vision encoder designed to recognize geometric premises from diverse styles of diagrams. To transfer the recognition to real-world PGPS benchmarks, we then propose a domain adaptation technique for GeoCLIP that leverages a small set of diagram–caption pairs from the target domains.
4.1 GeoCLIP
@@
For the caption of a diagram, we filter out some geometric properties from the original description of a diagram used to render the image. Specifically, we only keep the following four properties, concyclic, perpendicularity, angle measures, and length measures, from the visual premises shown in Table 5. After that, we convert the remaining descriptions written in the formal language into natural language. We filter out some properties for two reasons. First, some properties are not recognizable from the rendered diagram without additional information, e.g., congruency. These properties are listed as non-visual premises in Table 5. Second, collinearity and parallelity occur so frequently that they can marginalize others. Some examples of generated captions after filtering and translating are provided in the right-most column of Fig. 4. We call the filtered caption as GeoCLIP-style caption.
With this dataset, we fine-tune OpenCLIP (Radford et al., 2021) according to the CLIP objective which is formulated as:
@@
where $\mathcal{D}:=\{(D_{i},X_{i})\}_{i=1}^{N}$ is the diagram-caption pairs, $g$ is the vision encoder, $h$ is the text encoder, and $\tau$ is a temperature parameter. We named the resulting vision encoder as GeoCLIP. Section B.1 provides the details, including hyper-parameters.
We compare the performance of GeoCLIP to other self-supervised approaches trained with the same dataset. We test three self-supervised approaches: Jigsaw (Chen et al., 2021; Cao and Xiao, 2022), MAE (Ning et al., 2023; Xia et al., 2024), and VQ-VAE (Liang et al., 2023) used in previous work to improve the recognition performance of plane diagrams. We use the same architecture used for GeoCLIP for Jigsaw and MAE with the hyper-parameters used in the previous works. For VQ-VAE, we follow the architecture of Liang et al. (2023). All model performances are measured through the linear probing used in Section 3.2.
As shown in Table 1, GeoCLIP recognizes geometric features better than existing baselines and self-supervised methods. The self-supervised approaches generally perform poorly for the benchmark, justifying the choice of the objective. We also compare the performance of GeoCLIP against other encoders such as OpenCLIP. Note that although we outperform the other encoders in difficult tasks such as SquareShape and AngleDetection, these results might be unfair since the training set of GeoCLIP is similar to the diagrams in the benchmark. The t-SNE plots of the embeddings from the vision encoders are illustrated at Fig. 5.
We further ablate the filtering process in GeoCLIP. To this end, we compare GeoCLIP with its two variants: GeoCLIP (F $\times$), which uses the captions generated without filtering. We also test GeoCLIP (2K), which is trained on only 2,000 pairs, to see the effectiveness of the large-scale dataset. The results in Table 1 imply both the filtering and the training set size matter in enhancing geometric properties recognition.

### v2

@@ Abstract
We introduce GeoDANO, a geometric vision-language model (VLM) with a domain-agnostic vision encoder, for solving plane geometry problems. Although VLMs have been employed for solving geometry problems, their ability to recognize geometric features remains insufficiently analyzed. To address this gap, we propose a benchmark that evaluates the recognition of visual geometric features, including primitives such as dots and lines, and relations such as orthogonality. Our preliminary study shows that vision encoders often used in general-purpose VLMs, e.g., OpenCLIP, fail to detect these features and struggle to generalize across domains. To overcome the limitation, we develop GeoCLIP, a CLIP-based model trained on synthetic geometric diagram–caption pairs. Benchmark results show that GeoCLIP outperforms existing vision encoders in recognizing geometric features. We then propose our VLM, GeoDANO, which augments GeoCLIP with a domain adaptation strategy for unseen diagram styles. GeoDANO outperforms specialized methods for plane geometry problems and GPT-4o on MathVerse. The implementation is available at https://github.com/ml-postech/GeoDANO.
@@
Seunghyuk Cho1, Zhenyue Qin3, Yang Liu3, Youngbin Choi1, Seungbeom Lee1, Dongwoo Kim1,2 1Graduate School of Artificial Intelligence, POSTECH, 2Department of Computer Science and Engineering, POSTECH, 3Australian National University Correspondence to: Dongwoo Kim <dongwoo.kim@postech.ac.kr>
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in automated math problem solving, particularly through code-generation capabilities integrated with proof assistants (Moura and Ullrich, 2021; Nipkow et al., 2002; Chen et al., 2023; Wu et al., 2022; Hendrycks et al., 2021). Although LLMs excel at generating solution steps and correct answers in algebra and calculus (Zhou et al., 2024), their unimodal nature limits performance in plane geometry, where the solution depends on both diagram and text (Zhou et al., 2024).
Specialized vision-language models (VLMs) have accordingly been developed for plane geometry problem solving (PGPS) (Chen et al., 2021, 2022; Lu et al., 2021; Zhang et al., 2023; Zhang and Moshfeghi, 2024; Li et al., 2024b; Xia et al., 2024). Yet, whether these models genuinely leverage diagrams or rely almost exclusively on textual features remains unclear. This ambiguity arises because existing PGPS datasets typically embed sufficient geometric details within problem statements, potentially making the vision encoder unnecessary (Zhang and Moshfeghi, 2024). Fig.˜1 illustrates example questions from GeoQA and PGPS9K, where solutions can be derived without referencing the diagrams.
@@
Figure 1: Examples of diagram-caption pairs and their solution steps written in formal languages from the GeoQA and PGPS9k datasets. The problem description highlights the visual geometric premises and numerical variables in green and red, respectively. A significant difference in the style of the diagram and formal language can be observed.
We propose a new benchmark created via a synthetic data engine, which systematically evaluates the ability of VLM vision encoders to recognize geometric premises. Our empirical findings reveal that previously suggested self-supervised learning (SSL) approaches, e.g., vector quantized variational auto-encoder (VQ-VAE) (Liang et al., 2023) and masked auto-encoder (MAE) (Ning et al., 2023; Xia et al., 2024), and widely adopted encoders, e.g., OpenCLIP (Radford et al., 2021) and DinoV2 (Oquab et al., 2024), struggle to detect geometric features such as perpendicularity and degrees.
To this end, we propose GeoCLIP, a model pre-trained on a large corpus of synthetic diagram–caption pairs. By varying diagram styles (e.g., color, font size, resolution, line width), GeoCLIP learns robust geometric representations and outperforms prior SSL-based methods on our benchmark. Building on GeoCLIP, we introduce a few-shot domain adaptation technique that efficiently transfers the recognition ability to real-world diagrams. We finally propose GeoDANO by combining this domain-adapted GeoCLIP with an LLM, forming a domain-agnostic VLM for solving PGPS tasks.
In our experiments on MathVerse (Zhang et al., 2024a), which encompasses diverse plane geometry tasks and diagram styles, GeoDANO consistently outperforms both task-specific PGPS models and generalist VLMs. Ablation studies confirm the effectiveness of our domain adaptation strategy, showing improvements in optical character recognition (OCR)-based tasks and robust diagram embeddings across different styles.
We summarize the contributions as follows: We propose a novel benchmark for systematically assessing how well vision encoders recognize geometric premises in plane geometry diagrams (Section˜3); We introduce GeoCLIP, a vision encoder capable of accurately detecting visual geometric premises (Section˜4.1), and a few-shot domain adaptation technique that efficiently transfers this capability across different diagram styles (Section˜4.2); We show that our VLM, named GeoDANO, incorporating domain-adapted GeoCLIP, surpasses existing specialized PGPS VLMs and generalist VLMs on the MathVerse benchmark (Section˜5.2) and effectively interprets diverse diagram styles (Section˜5.3).
@@ 2 Related Work
In this section, we summarize the known PGPS benchmarks and models. Detailed comparison with previous work is reported in Appendix˜A.
@@ 2.1 PGPS benchmarks
Recently, MathVerse (Zhang et al., 2024a) provides an alternative view to existing PGPS benchmarks by directly encoding the geometric properties and relations into the diagrams rather than text description. Therefore, it is impossible to produce a valid solution without recognizing the necessary information from diagrams. CogAlign (Huang et al., 2025) introduces a benchmark that evaluates the spatial relationship understanding of pretrained vision encoders via linear probing. However, the questions in this benchmark focus purely on spatial relationships and do not involve symbols representing geometric relations.
@@ 2.2 Program generation based PGPS
Another line of research typically employs vision-language models (VLMs), where a VLM comprises a vision encoder and a language model (Zhang et al., 2023; Chen et al., 2021; Cao and Xiao, 2022; Ning et al., 2023; Chen et al., 2022; Liang et al., 2023; Xia et al., 2024; Li et al., 2024b). The vision encoder produces a visual embedding from the diagram, and the language model then generates solution steps in an autoregressive manner, conditioned on the textual description and the visual embedding. While the VLMs apply to various diagram formats, the visual geometric premises perception of the VLMs remains underexplored due to the abundance of textual information in existing benchmarks.
@@ 2.3 Contrastive learning in PGPS
Contrastive learning is applied in diverse domains such as computer vision (Schroff et al., 2015) and natural language processing (Gao et al., 2021). In the context of PGPS, contrastive learning is employed to address domain-specific challenges. GeoX (Xia et al., 2024) applies contrastive learning to the adapter layer of the VLM to enhance formal language comprehension. GeoGLIP (Zhang et al., 2025) utilizes grounded language-image pre-training with synthetic diagram and junction, boundary triples. Other approaches train the vision encoder itself using the contrastive language-image pre-training (CLIP) (Radford et al., 2021) objective: LANS (Li et al., 2024b) aligns patch embeddings from a vision Transformer (ViT) with text token embeddings if they describe the same point. MAVIS (Zhang et al., 2024b) employs diagram–caption pairs generated by a synthetic engine for CLIP, where the captions contain all the information in the diagram.
3 Benchmark for Geometric Premises
@@ 3.1 Benchmark preparation
  • Concyclic: A circle and four points are given. The task is identifying how many of those points lie on the circle.
  • TwoLines: Two lines, AB and BC, are given alongside other geometric objects. The task is determining whether AB and BC are perpendicular, collinear, or neither.
@@
An example of each task is provided in Fig.˜2.
@@
Our benchmark is built on top of AlphaGeometry (Trinh et al., 2024), which is designed to solve IMO-style plane geometry problems. The program provides useful functions such as formal language describing plane diagrams. The language predefines a set of geometric premises listed in Table˜A1, including all necessary properties to define our benchmark tasks. In addition, once a diagram description is given in formal language, the program renders a corresponding diagram with varying fonts, colors, widths, orientations, and resolutions, allowing us to have diagrams with diverse styles.
We create question-and-answer pairs based on AlphaGeometry. To sample a diverse set of questions and answers, we first establish a foundational geometric structure corresponding to the key problem of the task using the formal language provided by AlphaGeometry. For instance, in the AngleDetection benchmark, we specify AlphaGeometry problems that guarantee the presence of angle ABC with degree between 15° and 75° in the diagram.
We then execute Algorithm˜1 to generate diagrams and corresponding answers from these predefined geometric specifications. Specifically, Algorithm˜1 first samples an initial AlphaGeometry problem containing the essential geometric premise, then incrementally adds random geometric primitives and relations to diversify the diagrams. Finally, the answer to the generated diagram is determined from the initial sampled AlphaGeometry problem. Importantly, in the TwoLines, SquareShape, and AngleDetection benchmarks, the answer obtained from the initial AlphaGeometry problem remains unchanged and visually present despite adding other geometric elements. In contrast, for ObjectShape and Concyclic benchmarks, no additional geometric elements are introduced, further ensuring the accuracy of the answer. Consequently, the answers derived from these formal specifications consistently match the generated diagrams.
For each task, we generate 50,000, 10,000, and 10,000 question-and-answer pairs for training, validation, and testing, respectively. Additional details on the benchmark generation process are available at Section˜C.1.
@@ 3.2 Results
MAVIS-CLIP 91.64 71.78 58.84 60.48 16.12
GeoGLIP 98.40 91.97 60.22 63.10 11.58
AutoGeo 99.85 91.48 78.90 90.40 23.24
FM-ViT 96.89 87.29 61.73 64.25 15.90
SSL
Jigsaw 86.11 63.85 49.98 61.88 11.44
@@
GeoCLIP
GeoCLIP (F $\times$) 99.52 98.61 88.33 86.76 65.68
@@
We evaluate four eight adopted vision encoders for the open-sourced VLMs: OpenCLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), DinoV2 (Oquab et al., 2024), ConvNeXT (Liu et al., 2022), GeoGLIP (Zhang et al., 2025), FM-ViT (Lin et al., 2025), AutoGeo (Huang et al., 2024), and MAVIS-CLIP (Zhang et al., 2024b). For MAVIS-CLIP, FM-ViT, and AutoGeo, we train OpenCLIP on different datasets under their respective configurations: MAVIS-CLIP is trained on MAVIS-Caption following the MAVIS setup, AutoGeo is trained on synthetic diagram–caption pairs generated by AutoGeo using the MAVIS setup, and FM-ViT is trained on Geo170K (Gao et al., 2023) and MAVIS-Caption (Zhang et al., 2024b) following the FM-ViT setup.
To evaluate the vision encoder, we use linear probing, i.e., adding a linear layer on top of each encoder as a prediction head and training the linear layer from scratch while freezing the parameters of the vision encoder. We use a training set to train the prediction head and report the test accuracy with the best validation performance. The details for the hyper-parameters are described in Section˜C.2.
As shown in Table˜1, many existing vision encoders relatively well recognize the shape of objects but fail at the correct angle between two lines. The encoders also show some difficulties in recognizing the shape of a square and the relationship between two lines. Although the result may seem satisfactory at a glance, these errors will propagate to the downstream tasks when combined with LLMs.
4 GeoCLIP: Enhanced Vision Encoder
In this section, we first propose GeoCLIP, a new vision encoder designed to recognize geometric premises from diverse styles of diagrams. To transfer the recognition to real-world PGPS benchmarks, we then propose a domain adaptation technique for GeoCLIP that leverages a small set of diagram–caption pairs from target domains.
4.1 Training GeoCLIP
@@
For the caption of a diagram, we filter out some geometric properties from the original description of a diagram used to render the image. Specifically, we only keep the following four properties, concyclic, perpendicularity, angle measures, and length measures, from the visual premises shown in Table˜A1. After that, we convert the remaining descriptions written in the formal language into natural language. We filter out some properties for two reasons. First, some properties are not recognizable from the rendered diagram without additional information, e.g., congruency. These properties are listed as non-visual premises in Table˜A1. Second, collinearity and parallelity occur so frequently that they can marginalize others. Some examples of generated captions after filtering and translating are provided in the right-most column of Fig.˜A1. We call the filtered caption as GeoCLIP-style caption.
With this dataset, we fine-tune OpenCLIP (Radford et al., 2021) via the CLIP objective, formulated as:
@@
where $\mathcal{D}:=\{(D_{i},X_{i})\}_{i=1}^{N}$ is the diagram-caption pairs, $g$ is the vision encoder, $h$ is the text encoder, and $\tau$ is a temperature parameter. We named the resulting vision encoder as GeoCLIP. Section˜C.2 provides the details, including hyper-parameters.
We compare the performance of GeoCLIP to other self-supervised approaches trained with the same dataset. We test three self-supervised approaches: Jigsaw (Chen et al., 2021; Cao and Xiao, 2022), MAE (Ning et al., 2023; Xia et al., 2024), and VQ-VAE (Liang et al., 2023), used in previous work to improve the recognition performance of plane diagrams. We use the same architecture used for GeoCLIP for Jigsaw and MAE with the hyper-parameters used in the previous works. For VQ-VAE, we follow the architecture of Liang et al. (2023). All model performances are measured through the linear probing used in Section˜3.2.
As shown in Table˜1, GeoCLIP recognizes geometric features better than existing baselines and self-supervised methods. The self-supervised approaches generally perform poorly for the benchmark, justifying the choice of the objective. We also compare the performance of GeoCLIP against other encoders such as OpenCLIP. Note that although we outperform the other encoders in tasks such as SquareShape and AngleDetection, these results might be unfair compared to the existing pretrained vision encoders, since the training set of GeoCLIP is similar to the diagrams in the benchmark. The t-SNE plots of the embeddings from the vision encoders are illustrated at Fig.˜A3.
We further ablate the filtering process in GeoCLIP. To this end, we compare GeoCLIP with its two variants: GeoCLIP (F $\times$), which uses the captions generated without filtering. We also test GeoCLIP (2K), which is trained on only 2,000 pairs, to see the effectiveness of the large-scale dataset. The results in Table˜1 imply that both the filtering and the training set size matter in enhancing geometric properties recognition.